In a sample of 430 residents of Georgetown County, 165 reported they believed the county real estate taxes were too high.

Develop a 95 percent confidence interval for the proportion of residents who believe the tax rate is too high. (Round the p(proportion) value to 2 decimal places and final answers to 3 decimal places.)

  Confidence interval for the proportion of residents is  up to  .
Would it be reasonable to conclude that the majority of the taxpayers feel that the taxes are too high?
  
(Click to select)  Yes  No
Answer #1

\hat{p} =165/430 = 0.3837

For 95% confidence level , z_{c}=1.96

95% confidence interval for true proportion

\hat{p}\pm z_{c}*\sqrt{\hat{p}(1-\hat{p})/n}

= 0.3837 \pm 0.04596

= (0.338, 0.430)

No, it is not reasonable to conclude that majority of tax payers feel that taxes are too high.

As upper limit of confidence interval is less than 0.5 or 50% . So we can say that less than 50% feel that taxes are too high.